# Scaling Properties of Diffusion Models for Perceptual Tasks

### v2

@@ 1 Introduction
Diffusion models have emerged as powerful techniques for generating images and videos, while showing excellent scaling behaviors. However, diffusion models can also be utilized for downstream prediction for inverse vision problems. In this paper, we present a unified framework to perform depth estimation, optical flow estimation, and amodal segmentation with a single diffusion model, as illustrated in Figure 1.
Previous works such as Marigold (Ke et al., 2024), FlowDiffuser (Luo et al., 2024), and pix2gestalt (Ozguroglu et al., 2024) demonstrate the potential of repurposing image diffusion models for various inverse vision tasks individually. Building on this foundation, we perform an extensive empirical study, establishing scaling power laws for depth estimation, and display their transferability to other perceptual tasks. Using insights from these scaling laws, we formulate compute-optimal recipes for diffusion training and inference. We are the first to show that efficiently scaling compute for diffusion models leads to significant performance gains in downstream perceptual tasks.
Recent works have also focused on scaling test-time compute to enhance the capabilities of modern LLMs, as demonstrated by OpenAI’s o1 model (OpenAI, 2024). As Noam Brown noted, “It turned out that having a bot think for just 20 seconds in a hand of poker got the same boosting performance as scaling up the model by 100,000x and training it for 100,000 times longer.” In our experiments, we realize a similar trade off between allocating more compute during training versus test-time for diffusion models.

### v3

@@ 1 Introduction
Diffusion models have emerged as powerful techniques for generating images and videos, while showing excellent scaling behaviors. In this paper, we present a unified framework to perform a variety of perceptual tasks — depth estimation, optical flow estimation, and amodal segmentation — with a single diffusion model, as illustrated in Figure 1.
Previous works such as Marigold (Ke et al., 2024), FlowDiffuser (Luo et al., 2024), and pix2gestalt (Ozguroglu et al., 2024) demonstrate the potential of repurposing image diffusion models for various inverse vision tasks individually. Building on these prior works, we perform an extensive empirical study, establishing scaling power laws for depth estimation, and display their transferability to other perceptual tasks. Using insights from these scaling laws, we formulate compute-optimal recipes for diffusion training and inference. We find that efficiently scaling compute for diffusion models leads to significant performance gains in downstream perceptual tasks.
Recent works in other fields have also focused on scaling test-time compute to enhance the capabilities of modern LLMs, as demonstrated by OpenAI’s o1 model (OpenAI, 2024). Noam Brown, one of the key authors, expressed it quite pithily in a Ted Talk, “It turned out that having a bot think for just 20 seconds in a hand of poker got the same boosting performance as scaling up the model by 100,000x and training it for 100,000 times longer.” In our experiments, we observe a similar trade off between allocating more compute during training versus test-time for diffusion models with respect to downstream performance on perceptual tasks.